OpenStudy (anonymous):

I need someone to work me through this please!Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (agent0smith):

You'll have to try to learn to use the equation editor... don't make people decipher things like this "3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x" \[\Large 3\sqrt{2x} + x \sqrt {8x} - 5\sqrt{18x}\]

OpenStudy (anonymous):

I'm sorry, I wasn't thinking

OpenStudy (anonymous):

yes that's the equation and he got \[-10\sqrt{2x}\]

OpenStudy (agent0smith):

start simplifying it by looking for perfect squares\[\Large 3\sqrt{2x} + x \sqrt {4*2x} - 5\sqrt{9*2x}=\] \[\Large 3\sqrt{2x} + x \sqrt {4}*\sqrt{2x} - 5\sqrt{9}*\sqrt{2x}=\]

OpenStudy (anonymous):

what does * mean?

OpenStudy (anonymous):

I don't understand

OpenStudy (agent0smith):

* means multiply. All I've done is simplify it a bit, by breaking up some square roots. What part don't you understand?

OpenStudy (agent0smith):

\[\Large 3\sqrt{2x} + x \sqrt {8x} - 5\sqrt{18x} = \] \[\Large 3\sqrt{2x} + x \sqrt {4*2x} - 5\sqrt{9*2x}= \] \[\Large 3\sqrt{2x} + x \sqrt {4}*\sqrt{2x} - 5\sqrt{9}*\sqrt{2x}=\]

OpenStudy (anonymous):

do I multiply the inside with the outside? or do I find the square root of the inside and multiply that by the outside?

OpenStudy (anonymous):

3x+2x*sqrt 2x- 15* sqrt 2x?

OpenStudy (agent0smith):

Just simplify any pieces you can...like the sqrt of 4. And the sqrt of 9.

OpenStudy (agent0smith):

\[\Large 3\sqrt{2x} +2x \sqrt{2x} - 15\sqrt{2x}=\]

OpenStudy (anonymous):

then what do I do?

OpenStudy (agent0smith):

Now combine like terms. What would 3x+2x-15x be equal to?

OpenStudy (agent0smith):

Sorry, i corrected it. They all have the same srt2x, so treat it the same way.

OpenStudy (anonymous):

-11x sqrt 2x

OpenStudy (anonymous):

did I just solve it :D

OpenStudy (anonymous):

no its -9x!

OpenStudy (agent0smith):

Almost. Maybe try factoring out the common factor sqrt 2x \[\Large 3\sqrt{2x} +2x \sqrt{2x} - 15\sqrt{2x}= \] \[\Large \sqrt{ 2x} (3 +2x - 15)= \]

OpenStudy (anonymous):

do I add 12 to 2x?

OpenStudy (anonymous):

-12*

OpenStudy (agent0smith):

Which 12? you can change the order like so \[\Large \sqrt{ 2x} (3 -15 +2x)=\]

OpenStudy (agent0smith):

No, you can't combine 2x and -12, they are not like terms.

OpenStudy (anonymous):

so -12+2x sqrt 2x

OpenStudy (agent0smith):

\[\Large \sqrt{ 2x} (-12 +2x)=\] Now distribute the sqrt2x back into the brackets

OpenStudy (agent0smith):

Now you can see the guy's mistake... he combined -12 and 2x into -10x

OpenStudy (anonymous):

thank you! so the solution is?

OpenStudy (agent0smith):

Just distribute that sqrt2x back into the brackets

OpenStudy (anonymous):

hmm -12+2xsqrt2x?

OpenStudy (agent0smith):

You didn't distribute it correctly, it has to distribute to BOTH terms.

OpenStudy (anonymous):

im confused :(

OpenStudy (agent0smith):

A(B + C) = AB + AC Try using that on \[\Large \sqrt{ 2x} (-12 +2x)=\]

OpenStudy (anonymous):

-12x+2x 10x

OpenStudy (agent0smith):

Wait... no. Make sure you don't change anything like that. \[\Large \sqrt{ 2x} (-12 +2x)= -12 \sqrt{2x} + 2x \sqrt {2x}\]

OpenStudy (anonymous):

ohh ok sorry

OpenStudy (anonymous):

so that's the answer

OpenStudy (agent0smith):

Yes. Just be careful using the distributive property.

OpenStudy (agent0smith):

You can also do it this way (might've been easier). Change the order: \[\Large 3\sqrt{2x} +2x \sqrt{2x} - 15\sqrt{2x}= \] \[ \Large 3\sqrt{2x} - 15\sqrt{2x} +2x \sqrt{2x} =\]

OpenStudy (anonymous):

alright thank you!!

OpenStudy (agent0smith):

Welcome :)